### v2

# Evolving Language Models without Labels: Majority Drives Selection, Novelty Promotes Variation

Abstract
Large language models (LLMs) are increasingly trained with reinforcement learning from verifiable rewards (RLVR), yet real-world deployment demands models that can self-improve without labels or external judges. Existing self-improvement approaches primarily rely on self-confirmation signals (e.g., confidence, entropy, or consistency) to generate rewards. This reliance drives models toward over-confident, majority-favored solutions, causing an entropy collapse that degrades pass@$n$ and reasoning complexity. To address this, we propose Evol-RL, a label-free framework that mirrors the evolutionary principle of balancing selection with variation. Concretely, Evol-RL retains the majority-voted answer as an anchor for stability, but adds a novelty-aware reward that scores each sampled solution by how different its reasoning is from other concurrently generated responses. This majority-for-stability + novelty-for-exploration rule mirrors the variation–selection principle: selection prevents drift, while novelty prevents collapse. Evaluation results show that Evol-RL consistently outperforms the majority-only baseline; e.g., training on label-free AIME24 lifts Qwen3-4B-Base AIME25 pass@1 from baseline’s 4.6% to 16.4%, and pass@16 from 18.5% to 37.9%. Evol-RL not only prevents in-domain diversity collapse but also improves out-of-domain generalization (from math reasoning to broader tasks, e.g., GPQA, MMLU-Pro, and BBEH).
[Uncaptioned image] Code             [Uncaptioned image] Models
1 Introduction
The reasoning capabilities of Large Language Models (LLMs) have advanced dramatically, particularly through paradigms like Reinforcement Learning with Verifiable Rewards (RLVR) (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025). The next frontier of intelligence lies in enabling LLMs to autonomously evolve, continuously learning from the vast, unlabeled data streams they encounter in real-world environments. This label-free evolving paradigm allows a model to iteratively improve itself while solving tasks, without relying on ground-truth labels or external judges, making it both practical and necessary. However, turning inference into learning reopens a long-standing RL problem: balancing exploration and exploitation. This dilemma becomes especially severe in label-free settings, where models must rely on internal signals (e.g., inherent self-consistency, entropy, or confidence) to generate rewards for themselves (Grandvalet & Bengio, 2004; Lee et al., 2013; Zuo et al., 2025; Shafayat et al., 2025; Li et al., 2025b).
Figure 1: TTRL’s entropy collapse vs. EVOL-RL’s diversity preservation on Qwen3-4B-Base (trained label-free on MATH-500). Majority-only TTRL drives pass@$n>1$ down, shortens reasoning, and collapses entropy, whereas Evol-RL improves accuracy, sustains reasoning diversity.
The fundamental flaw in relying on internal signals is not merely that they are initially noisy or biased, but that the learning process itself actively degrades the quality of the reward signal over time (Liang et al., 2024; Zeng et al., 2024). By rewarding conformity to its self-confirmation, the model systematically eliminates the solution diversity (Lee et al., 2024). This creates a degenerative feedback loop: a progressively narrower and more biased policy generates an increasingly impoverished reward signal, which in turn accelerates the policy’s collapse into a low-entropy state (Ding et al., 2025; Liang et al., 2024). Similar dynamics are well known in RL and self-training when entropy regularization or external supervision is absent (Haarnoja et al., 2018). Recent studies also show that training on self-generated data can harm diversity over time (Shumailov et al., 2024) and eventually lead to collapse. Figure 1 illustrates this phenomenon in reasoning: under traditional Test-Time Reinforcement Learning (TTRL) (Zuo et al., 2025), pass@1 may rise but pass@$n$ drops, while response length and complexity steadily decline, indicating that the model fails to evolve.
In this paper, we ground LLM evolving in the simple rule behind biological evolution: variation creates new candidates; selection keeps what works. Existing methods effectively implement only the selection half of evolution, driving the population toward whatever the model already believes. This majority-only (or entropy minimization, confidence maximization) reinforcement amplifies existing biases and often leads entropy collapse and shrinking response diversity, as shown above. Our formulation restores the full evolutionary loop: we pair selection, which stabilizes optimization by keeping high-quality solutions, with variation, which explicitly promotes novelty and sustains exploration. This idea is deeply rooted in decades of evolutionary computation research, including genetic algorithms (Holland, 1992; Eiben & Smith, 2015), novelty search (Lehman & Stanley, 2011), and quality–diversity (QD) methods such as MAP-Elites (Pugh et al., 2016), which collectively show that diversity preservation is essential for avoiding collapse and enabling robust, long-term progress.
Hence, we propose EVolution-Oriented and Label-free Reinforcement Learning (Evol-RL), a simple objective that combines a stabilizing selection signal with an explicit variation incentive. Concretely, Evol-RL retains the majority-voted answer as the anchor for stability, but adds a novelty-aware reward that scores each sampled solution by how different its reasoning is from other concurrently generated responses (semantic similarity of their reasoning traces). This majority-for-stability + novelty-for-exploration rule mirrors the variation–selection principle: selection prevents drift; novelty prevents collapse. As demonstrated in Figure 1, Evol-RL successfully averts all symptoms of diversity collapse, fostering a healthy equilibrium between refining known solutions and discovering new ones. This balanced approach translates into substantial performance gains, especially in out-of-domain generalization. For instance, after training on AIME24, Evol-RL elevates the Qwen3-4B-base model’s pass@1 accuracy on the AIME25 benchmark from 4.6% (TTRL) to 16.4%, while more than doubling the pass@16 accuracy from 18.5% to 37.9%.
Contributions. (1) We diagnose why majority-only objectives shrink exploration during label-free training and formalize their link to entropy collapse on reasoning tasks. (2) We provide a new perspective on label-free learning by framing it as an evolutionary system. This view allows us to diagnose diversity collapse as a form of premature convergence and solve it by applying the core evolutionary principle of balancing selection with variation. (3) We design a practical novelty-aware reward that complements majority selection and enables stable, label-free improvement. Across math benchmarks, Evol-RL reverses the pass@$n$ decline, maintains longer and more informative chains of thought, and improves out-of-domain accuracy, while remaining simple to implement. (4) We deliver state-of-the-art results in unsupervised RL training, demonstrating that Evol-RL achieves significant out-of-domain generalization gains where prior methods fail, such as more than tripling pass@1 accuracy and doubling pass@16 accuracy on AIME25 benchmark.
2 Related Work
Enhancing Reasoning in Large Language Models.
Significant progress in LLM reasoning has been driven by RLVR (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025; Yu et al., 2025; Xiong et al., 2025; Dai et al., 2025b), which fine-tunes models using RL on tasks where an automated verifier can confirm the correctness of the final answer, such as mathematics and coding (Zeng et al., 2025; Wang et al., 2025a; b; Cui et al., 2025; Huang et al., 2025; Dai et al., 2025a; Zheng et al., 2025b; Zhou et al., 2025b; Zheng et al., 2025a; Fang et al., 2025). While highly effective, the reliance of RLVR on external verifiers restricts its applicability to domains with deterministic, easily checkable solutions (Zhao et al., 2025b; Zhou et al., 2025a; 2024). Our work contributes to the effort of improving reasoning in more general domains where such verifiers are unavailable.
Label-Free Adaptation and Self-Improvement.
To overcome the limitations of verifiers and adapt to new data distributions, researchers have focused on label-free learning methods that generate reward signals without ground-truth labels. These approaches primarily fall into two categories. One line of research derives rewards from the model’s intrinsic confidence, training the model to become more "certain" by rewarding low-entropy or self-consistent outputs (Prabhudesai et al., 2025; Agarwal et al., 2025; Zhao et al., 2025a; Zhang et al., 2025b; Shafayat et al., 2025; Chung et al., 2025; Li et al., 2025a; Zhang et al., 2025a). The other prominent paradigm, which our work directly addresses, bootstraps supervision from majority. Test-Time Reinforcement Learning (TTRL) exemplifies this by using the majority-voted answer from multiple samples as a pseudo-label for RL updates (Zuo et al., 2025). While empirically powerful, we identify a critical flaw in the majority-driven approach: it suppresses solution diversity and actively punishes correct but non-mainstream reasoning, leading to the entropy collapse we describe. While ETTRL adjusts exploration within the original self-consistency framing (Liu et al., 2025), we are the first to pin down the majority trap and redesign the learning target to couple majority with population-level diversity.
3 Method
Figure 2: An overview of the Evol-RL framework. For each prompt, the policy generates multiple responses. These are grouped by their final answer to identify the majority group. A novelty score is then computed for each response based on its semantic dissimilarity to others. Finally, a reward is assigned based on both majority (selection) and novelty (variation), guiding the policy update via GRPO. In the illustration, colors group responses by their final answer, while different marker shapes indicate semantically distinct reasoning paths.
Our approach is illustrated in Figure 2. which uses Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as its optimization algorithm, but guides it with a novel reward function that explicitly balances majority with novelty.
3.1 Optimization with GRPO
GRPO is a policy-gradient algorithm designed for fine-tuning LLMs without a separate value function. Its central idea is to evaluate each sampled response relative to a group of its peers generated for the same prompt. This relative evaluation is then used to update the policy with a PPO-style clipped objective, regularized by a KL penalty to ensure stable learning.
For a given prompt $\mathbf{q}$, a policy LLM $\pi_{\theta_{\text{old}}}$ generates a group of $G$ complete responses $\{\mathbf{o}_{1},\ldots,\mathbf{o}_{G}\}$. Each response $\mathbf{o}_{i}$ receives a scalar reward $r_{i}$. Rewards within the group are normalized with a z-score to obtain a response-level advantage:
$\hat{A}_{i}=\frac{r_{i}-\text{mean}(r_{1},\ldots,r_{G})}{\text{std}(r_{1},\ldots,r_{G})},$
The policy is optimized with a clipped surrogate objective:
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\!\left\{\frac{\pi_{\theta}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)}{\pi_{\theta_{\mathrm{old}}}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)}\,\hat{A}_{i,t},\;\operatorname{clip}\!\left(\frac{\pi_{\theta}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)}{\pi_{\theta_{\mathrm{old}}}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)},\,1-\epsilon_{\mathrm{low}},\,1+\epsilon_{\mathrm{high}}\right)\hat{A}_{i,t}\right\}$ (1)
3.2 Reward Design: Implementing Selection and Variation
Our reward design directly implements the principles of selection and variation to counteract diversity collapse. Selection, based on correctness via majority vote, provides a stable signal to prevent the policy from drifting. Variation, driven by semantic novelty, provides the exploratory pressure needed to maintain a diverse set of reasoning strategies.
A key design choice is that the novelty incentive is applied strategically to all solutions—both those that agree with the majority and those that do not. For majority-aligned solutions, rewarding novelty encourages the model to discover multiple valid reasoning paths to the correct answer, directly fighting the decline in pass@n performance. For minority solutions, rewarding novelty is crucial for escaping local optima. It discourages policy collapse into a few high-frequency failure modes and instead incentivizes exploration of the broader reasoning space, which is essential for increasing the probability of discovering a previously inaccessible, correct solution path. This integration transforms the learning process: it not only mitigates diversity collapse in the current task but also aligns with the goals of continual learning. By preserving multiple reasoning modes while anchoring to a correct solution, Evol-RL avoids forgetting potentially useful strategies and retains knowledge diversity for future tasks. Thus, training under Evol-RL becomes not only an optimization for present performance but also a proactive investment in future adaptability.
Reward Formulation.
For each prompt, the policy samples $G$ responses $\{o_{i}\}_{i=1}^{G}$. Each response is scored on three criteria:
[…]
2. Majority (Selection): A binary label $y_{i}\in\{+1,-1\}$ is assigned based on whether a response’s answer matches the majority-voted answer from the valid responses. This serves as our selection signal.
3. Novelty (Variation): We compute embeddings for the reasoning part of each response to form a cosine similarity matrix. For each response $o_{i}$, we calculate its mean similarity $\bar{s}_{i}$ to other responses in the same group (i.e., either majority or minority) and its maximum similarity $m_{i}$ to any other response in the entire batch. The mean similarity is calculated on an intra-group basis because the majority and minority solutions are often semantically distant; a global mean would be dominated by this gap, obscuring the finer-grained variations among peer solutions within the majority group. The novelty score is:
$u_{i}\;=\;1-\big(\alpha\,\bar{s}_{i}+(1-\alpha)\,m_{i}\big),\qquad\alpha\in~(\text{default }0.5).$
This score is designed to penalize two distinct forms of redundancy: a high $\bar{s}_{i}$ indicates conformity to the group’s semantic average, while a high $m_{i}$ flags near-duplication of another specific response. The score promotes both local and global diversity. Finally, we min-max normalize the scores $\{u_{i}\}$ separately within the majority and minority groups to get $\tilde{u}_{i}$. This intra-group normalization is crucial, as it ensures that novelty is measured relative to one’s direct peers, allowing for a fair comparison of diversity within each group.
Final Reward Mapping.
We map the majority label and normalized novelty score into non-overlapping reward bands. This ensures that the selection signal from the majority vote is always prioritized, while novelty refines the reward within each group:
[…]
Critically, this structure guarantees that any majority solution, regardless of its novelty, receives a higher reward than any minority solution. This maintains a strong pressure towards correctness. More details about the reward implementation are presented in Appendix A.4
Supporting Mechanisms.
To further reinforce this reward design, we employ two complementary mechanisms. First, within the GRPO objective (Eq. 1), we use an asymmetric clipping range ($\epsilon_{\text{high}}>\epsilon_{\text{low}}$) (Yu et al., 2025). This allows promising and novel solutions with high advantages to receive larger gradient updates, preventing them from being prematurely clipped. Second, we add a token-level entropy regularizer to maintain diversity during the initial generation process:
$\mathcal{L}_{\mathrm{ent}}(\theta)=-\lambda_{\mathrm{ent}}\;\mathbb{E}_{o\sim\pi_{\theta}}\left[\frac{1}{|o|}\sum_{t=1}^{|o|}H\!\left(\pi_{\theta}(\cdot\mid o_{<t},x)\right)\right],\qquad H(p)=-\sum_{v}p(v)\log p(v).$ (2)
The total objective, $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{GRPO}}+\mathcal{L}_{\text{ent}}$, thus directs learning toward semantically distinct, high-quality responses while maintaining a diverse population of solutions.
3.3 How EVOL-RL Avoids Collapse Through an Evolutionary Analogy.
EVOL-RL avoids this failure mode by mirroring biological evolution, which balances a stabilizing Selection pressure with a dynamic Variation mechanism. The majority vote acts as our Selection pressure, providing a crucial anchor to correctness. By itself, however, this would lead to a uniform population of solutions, vulnerable to collapse, much like a species with no genetic diversity is vulnerable to a single disease.
To prevent this, our three-part Variation strategy creates a robust exploratory dynamic. The entropy regularizer acts like a higher "mutation rate," ensuring a constant supply of diverse solutions for the system to work with. The novelty reward then provides directional pressure to this variation, giving a "survival bonus" to solutions that are semantically different from their peers. Finally, asymmetric clipping ensures that when a highly beneficial "mutation"—a rare, novel, and correct solution—appears, its strong learning signal is fully preserved for the next generation.
This design makes a collapsed state inherently unstable. In a uniform population of near-duplicate solutions, any "mutation" is, by definition, highly novel and receives a higher reward. The learning algorithm is thus forced to shift probability away from the uniform cluster and toward this more promising, distinct solution, ensuring that the policy remains robustly diverse.
4 Experiments
4.1 Experimental Setup
Benchmarks. To test our method at scale, we use the large, standard MATH training set (MATH-TRAIN) (Hendrycks et al., 2021). We also follow the TTRL (Zuo et al., 2025) by training on two much smaller test sets: the general-purpose MATH-500 and the competition-level AIME24 (Li et al., 2024). This comprehensive setup allows us to validate Evol-RL’s versatility across both large-scale and specialized training conditions. Critically, during all training runs, we use only the problem statements, without any ground-truth labels or solutions. For evaluation, we assess the performance of our trained models on a diverse set of five benchmarks to measure both in-domain and out-of-domain generalization. The evaluation suite includes AIME24, AIME25, MATH500, AMC (Li et al., 2024), and GPQA-Diamond (GPQA) (Rein et al., 2024). Detailed training configuration can be found in Appendix A.
4.2 Main Results
Table 1: Comparison of models trained with TTRL and Evol-RL. Each cell shows pass@1/pass@16 (averaged on 32 rollouts). $\Delta$ uses red (+) for positive and blue for negative values, showing the difference between w/Evol-RL and w/TTRL.
Training Dataset Model MATH AIME24 AIME25 AMC GPQA
Qwen3-4B-Base
Base Model 67.4/89.6 10.0/32.4 5.5/30.0 39.3/75.2 34.4/85.6
MATH-TRAIN w/TTRL 75.4/86.9 12.1/23.2 6.8/28.6 42.5/75.2 36.5/81.4
w/Evol-RL 80.0/93.3 20.7/47.6 17.5/39.9 51.4/80.3 37.2/88.7
$\Delta$ +4.6/+6.4 +8.6/+24.4 +10.7/+11.3 +8.9/+5.1 +0.7/+7.3
MATH-500 w/TTRL 79.3/83.2 10.0/28.0 7.2/29.9 47.6/72.0 36.2/75.9
w/Evol-RL 79.8/93.8 19.0/43.2 16.1/41.9 50.3/82.2 38.8/89.1
$\Delta$ +0.5/+10.6 +9.0/+15.2 +8.9/+12.0 +2.7/+10.2 +2.6/+13.2
AIME24 w/TTRL 73.8/84.5 16.7/16.7 4.6/18.5 43.6/65.8 35.1/73.5
w/Evol-RL 79.6/93.6 20.6/40.9 17.1/42.0 49.9/80.9 38.0/87.8
$\Delta$ +5.8/+9.1 +3.9/+24.2 +12.5/+23.5 +6.3/+15.1 +2.9/+14.3
Qwen3-8B-Base
Base Model 63.6/91.5 12.0/39.4 8.2/30.8 38.7/77.6 34.9/88.0
MATH-TRAIN w/TTRL 81.1/91.1 16.7/37.6 15.6/35.9 53.6/74.0 38.1/77.1
w/Evol-RL 83.6/94.1 26.0/51.7 21.6/43.1 55.5/86.1 43.5/88.1
$\Delta$ +2.5/+3.0 +9.3/+14.1 +6.0/+7.2 +1.9/+12.1 +5.4/+11.0
MATH-500 w/TTRL 85.7/91.9 17.7/40.1 16.5/34.3 51.1/79.1 43.5/84.0
w/Evol-RL 84.7/95.1 24.1/49.5 20.2/44.4 58.8/86.0 43.9/92.2
$\Delta$ -1.0/+3.2 +6.4/+9.4 +3.7/+10.1 +7.7/+6.9 +0.4/+8.2
AIME24 w/TTRL 76.8/86.2 20.0/20.0 11.4/25.4 49.5/69.1 38.3/74.7
w/Evol-RL 83.1/94.2 25.4/38.1 16.5/34.7 54.4/85.8 45.2/90.0
$\Delta$ +6.3/+8.0 +5.4/+18.1 +5.1/+9.3 +4.9/+16.7 +6.9/+15.3
The main results of our experiments are summarized in Table 1. We highlight four key findings that demonstrate the superiority of Evol-RL over the majority-only TTRL baseline.
Evol-RL Enhances Both Pass@1 and Pass@16 Performance. Across all experimental settings, Evol-RL consistently and substantially improves ‘pass@16‘ performance over TTRL, with gains frequently exceeding 20 percentage points on the most challenging benchmarks (e.g., +24.2pp on AIME24 for the 4B model). Evol-RL also delivers more consistent and substantial improvements to pass@1 accuracy than TTRL. This demonstrates that our method strengthens not only the model’s single-shot accuracy but also its ability to explore through multiple attempts.
Consistent Improvements Across Model Scales and Training Data Sizes. The benefits of Evol-RL are robust across both the 4B and 8B model scales, and critically, across training datasets of vastly different sizes. The performance improvements hold true whether training on the large-scale MATH-TRAIN set or the smaller, more specialized MATH-500 and AIME24 sets. This suggests that the underlying mechanism—balancing a majority anchor with novelty-driven rewards—is a fundamental improvement that scales effectively with both model capacity and data volume.
Strong Cross-Difficulty Robustness Evol-RL demonstrates powerful generalization, learning abstract reasoning skills that transfer effectively across different mathematical domains. A powerful example is seen with the 4B model: when trained exclusively on MATH-500, its pass@16 performance on AIME24 and AIME25 is nearly identical to the performance achieved when training on AIME24 directly, confirming that Evol-RL learns fundamental skills rather than simply overfitting. This effect is further amplified by scale; for the 8B model, training on the large MATH-TRAIN dataset yields pass@1 performance on AIME24 (26.0%) and AIME25 (21.6%) that is far superior to training on AIME24 directly (25.4% and 16.5% respectively). This indicates that Evol-RL effectively leverages both specialized and large-scale data to build fundamental and transferable reasoning abilities.
Generalization on Non-Mathematical Reasoning Tasks. The advantages of Evol-RL extend beyond the domain of mathematics. On the GPQA benchmark, where TTRL consistently causes pass@16 performance to degrade compared to the base model, Evol-RL reliably recovers and surpasses the base model. Across all training configurations, it achieves gains of +7 to +15 pp in pass@16 over TTRL, indicating that our diversity-preserving reward mechanism fosters a more generalizable reasoning ability that transfers effectively across different domains.
4.3 Ablation Study
Table 2: Performance of Qwen3-4B-Base with Evol-RL and its ablations on five benchmarks. Each cell reports pass@1/pass@16 accuracy.
Training Dataset Model MATH AIME24 AIME25 AMC GPQA
MATH-500 w/Evol-RL 79.8/93.8 19.0/43.2 16.1/41.9 50.3/82.2 38.8/89.1
-ClipHigh 75.1/91.8 12.2/31.8 11.4/31.3 42.7/73.9 32.3/81.8
-Ent 79.5/93.4 18.3/38.5 14.7/34.3 48.3/78.6 38.6/87.0
-ClipHigh-Ent 76.3/92.6 12.8/38.8 12.5/37.4 46.2/77.4 35.6/88.8
-Novelty Reward 79.3/88.7 12.1/27.0 11.1/34.8 47.6/73.3 37.9/81.4
AIME24 w/Evol-RL 79.6/93.6 20.6/40.9 17.1/42.0 49.9/80.9 38.0/87.8
-ClipHigh 74.1/89.4 14.1/26.7 8.1/31.1 44.6/73.2 35.3/81.5
-Ent 66.7/89.8 10.0/31.4 6.6/27.8 38.7/74.2 34.0/86.2
-ClipHigh-Ent 75.3/89.0 16.6/26.9 9.2/32.2 45.8/71.2 37.1/82.0
-Novelty Reward 79.4/93.0 17.7/35.6 15.9/37.4 48.8/79.6 37.9/87.1
Setup. We conduct an ablation study on Evol-RL-trained models on Qwen3-4B-Base. Evol-RL introduces three key modifications compared to the TTRL baseline: (i) the novelty-aware reward function, (ii) a rollout entropy regularizer to encourage exploration, and (iii) an asymmetric PPO clipping window (higher "ClipHigh") to better preserve learning signals from high-reward samples. We systematically remove these components one at a time ("-Novelty Reward", "-Ent", "-ClipHigh") or in combination. The results are reported in Table 2.
The Critical Role of Novelty on Easier Datasets. The importance of the novelty reward is most evident when the model is trained on the MATH-500 dataset. Removing it causes the largest performance degradation in pass@16, especially on the more difficult, out-of-domain AIME24/25. This is because on a dataset with lower complexity, a majority-only approach can quickly cause the model to lock into a single, repetitive reasoning template. Our novelty reward prevents this template lock-in and promotes generalizable skills.
Exploration Mechanisms as Critical Enablers on Harder Tasks. On more challenging datasets like AIME24, where the inherent problem difficulty naturally induces a higher baseline of exploration, the other two components become more critical. In this setting, removing the entropy regularizer or the asymmetric clipping consistently lowers pass@16 performance on AIME-style problems. These mechanisms act as crucial enablers for the novelty reward: the entropy regularizer ensures a rich and continuous supply of varied reasoning paths for the novelty selector to act upon, while the higher clipping threshold preserves the full learning signal from rare but high-value solutions.
4.4 Training Dynamics: How Evol-RL Escapes Entropy Collapse
Figure 3: Training dynamics for Evol-RL and TTRL. Left: models trained on MATH-TRAIN. Middle: models trained on MATH-500. Right: models trained on AIME24. Each panel plots, over training steps, (i) Pass@1 on AIME25, (ii) average response length on the training set, and (iii) policy entropy on the training set.
To understand the reasons for Evol-RL’s better performance, we analyze its training dynamics in comparison to TTRL in a label-free setting, as shown in Figure 3. An analysis of the training dynamics for the 8B models is presented in Appendix B.2.
Stage 1: Initial Collapse Under Majority Signal. Across all three training settings, a consistent initial dynamic unfolds: both Evol-RL and TTRL show a sharp drop in policy entropy and average response length. This initial phase demonstrates the powerful homogenizing effect of the majority-driven reward, which quickly pushes both models toward short, high-frequency response templates. For TTRL, this collapsed state proves to be permanent; it remains trapped in this low-entropy, low-complexity state for the duration of the training run, regardless of the dataset’s scale or difficulty.
Stage 2: The Evolving Point and Coordinated Recovery. Following the initial collapse, the training dynamics reveal a crucial divergence centered around a distinct "evolving point". Before this point, Evol-RL’s trajectory is nearly indistinguishable from TTRL’s; both models exhibit similar performance values and trends, dominated by the majority signal. However, a clear inflection point consistently emerges for Evol-RL, after which its performance rapidly improves. While the exact timing of this "evolving point" varies across datasets, its appearance is a robust feature of our method. After this "evolving point", Evol-RL enters a recovery phase characterized by a sustained and coordinated rise across all key metrics: policy entropy breaks away from near-zero values, average response length increases, and out-of-domain accuracy steadily climbs. This coordinated recovery allows the model to reach a new, significantly higher performance plateau where it eventually stabilizes, demonstrating its ability to break free from the majority trap.
Evol-RL’s ability to escape the collapsed state comes from the synergy of its three core components. The entropy regularizer ensures a continuous supply of diverse rollouts, preventing the initial search space from becoming completely uniform. The asymmetric clipping preserves the full gradient signal from the rare but high-value "majority-and-novel" samples that are crucial in the early training phase. Finally, the novelty reward acts as a selection pressure, consistently re-ranking credit within the majority group to favor these distinct solutions over their near-duplicate peers.
4.5 Evol-RL Components Also Strengthen Supervised GRPO (RLVR)
Figure 4: Performance of Evol-RL’s exploration-enhancing components when applied to a standard supervised GRPO baseline. The Qwen3-4B-Base model is trained on the MATH trainig set (Hendrycks et al., 2021) with a ground-truth verifier (RLVR).
Setup. We apply Evol-RL’s three exploration-enhancing ingredients to a standard supervised GRPO baseline trained on MATH training set (Hendrycks et al., 2021) with a ground-truth verifier (RLVR) for two epochs. Figure 4 reports the results.
The primary finding is that the three components are still synergistic, with their full combination yielding the most significant and consistent performance improvements. This complete configuration, GRPO+ClipHigh+Ent+Novelty, boosts pass@16 accuracy by 7% to 12% on the challenging out-of-domain AIME24 and AIME25 benchmarks. Crucially, these gains are achieved while also improving pass@1 accuracy, demonstrating that the mechanisms enhance multi-path reliability without sacrificing single-shot performance. This robust improvement extends across all evaluation benchmarks, including the cross-domain GPQA task, demonstrating the great potential of variation reward in a broader context.
Table 3: Generalization performance of the Qwen3-8B-Base model on broader reasoning benchmarks after label-free training on MATH-TRAIN.
Model MMLU-Pro SuperGPQA BBEH
Pass@1 Pass@4 Pass@1 Pass@4 Pass@1 Pass@4
Qwen3-8B-Base 47.3 74.5 26.5 54.1 10.4 24.0
        w/TTRL 53.4 73.9 29.7 53.3 12.1 24.1
        w/EVOL-RL 55.3 78.5 30.2 57.0 11.5 24.9
4.6 Generalization to Broader Reasoning Benchmarks
To assess whether the reasoning skills enhanced by our method on mathematical data are fundamental and transferable, we evaluate our models on a suite of broader, non-mathematical reasoning benchmarks. After training the Qwen3-8B-Base model on the MATH-TRAIN dataset in a label-free setting, we measure its performance on MMLU-Pro (Wang et al., 2024), SuperGPQA (Team et al., 2025), and BBEH (Kazemi et al., 2025). The results, presented in Table 3, demonstrate that EVOL-RL fosters a more generalizable reasoning ability compared to TTRL.
A contrasting pattern emerges between the two methods. While TTRL shows clear improvements over the base model on pass@1 accuracy, its effect on pass@4 is less consistent, falling slightly below the base model’s performance on SuperGPQA and BBEH. This pattern is consistent with our findings on the mathematical reasoning tasks, where the narrow focus of the consensus-only objective can hurt multi-path reliability. In contrast, EVOL-RL demonstrates a more robustly positive transfer of skills, improving upon both the base model and TTRL across pass@1 and pass@4 metrics. For example, on MMLU-Pro, EVOL-RL achieves a pass@4 score of 78.5%, a clear improvement over TTRL’s 73.9%. This indicates that our principle of encouraging diverse reasoning helps the model learn more fundamental skills that generalize effectively beyond mathematics.
5 Conclusion
In this work, we diagnose the entropy collapse, a critical failure mode in LLM evolving where majority-only rewards suppress solution diversity and harm generalization. To solve this, we propose Evol-RL, a framework that balances the stability of majority-vote selection with an explicit variation incentive that rewards semantic novelty. Our experiments demonstrate that Evol-RL successfully prevents collapse by maintaining policy entropy and reasoning complexity, which translates into substantial performance gains on both in-domain and out-of-domain benchmarks. By anchoring learning to a stable majority signal while simultaneously encouraging exploration, Evol-RL offers a robust and practical methodology for enabling LLMs to continuously and autonomously evolve without external labels.
Appendix A Implementation Details
This section provides additional details on the implementation of our reward formulation and supporting mechanisms.
A.1 Training Configuration.
We conduct our experiments on two recent open-source base models: Qwen3-4B-Base and Qwen3-8B-Base. Our training process is implemented using the GRPO algorithm. We adopt a setup similar to that of TTRL for generating training signals. For each problem instance, we first perform a rollout phase where the policy generates $64$ candidate responses. A majority label is then determined by performing a majority vote on the final answers extracted from these $64$ samples. Subsequently, a random subset of $32$ of these responses is used to form a batch for a single model update step. To ensure that the model has sufficient capacity for complex, multi-step reasoning, we set the maximum response length to 12,288 tokens during generation. To guide the model’s reasoning process, we utilize the system prompt from SimpleRL-Zoo (Zeng et al., 2025). Implementation details are discussed in Appendix A.
A.2 System Prompt
For all experiments, we used the following system prompt to guide the model’s generation format, ensuring that it produces a step-by-step reasoning process and a clearly marked final answer (Zeng et al., 2025):
[…]
A.3 Answer and Reasoning Extraction
To implement the scoring criteria described in the main text, we apply the following extraction procedure for each generated response $o_{i}$:
[…]
A.4 Novelty Score Calculation Details
The novelty score $u_{i}$ relies on computing semantic similarity between the reasoning parts of the generated responses.
Embedding Model.
We use the Qwen3-4B-Embedding model to generate dense vector representations for the extracted reasoning parts. Each vector is L2-normalized before similarity computation.
Cosine Similarity Matrix.
For a group of $G$ responses with corresponding L2-normalized embedding vectors $\{\mathbf{v}_{1},\ldots,\mathbf{v}_{G}\}$, the cosine similarity matrix $\mathbf{S}\in\mathbb{R}^{G\times G}$ is computed as $\mathbf{S}=\mathbf{V}\mathbf{V}^{T}$, where $\mathbf{V}$ is the matrix whose rows are the vectors $\mathbf{v}_{i}$. The element $S_{ij}$ represents the cosine similarity between the reasoning of response $o_{i}$ and $o_{j}$.
Intra-Group Min-Max Normalization.
To obtain the normalized novelty score $\tilde{u}_{i}\in[0,1]$ from the raw scores $\{u_{k}\}$ within a specific group (e.g., the majority group), we apply standard min-max normalization:
$\tilde{u}_{i}=\frac{u_{i}-\min(\{u_{k}\})}{\max(\{u_{k}\})-\min(\{u_{k}\})+\epsilon_{\text{norm}}}$
where $\epsilon_{\text{norm}}$ is a small constant (e.g., $10^{-8}$) to prevent division by zero in cases where all novelty scores in the group are identical.
A.5 Hyperparameter Settings
For our label-free experiments, we largely follow the settings established by TTRL to ensure a fair comparison. The general hyperparameters are detailed in Table 4, and the settings specific to our Evol-RL method are listed in Table 5.
Table 4: General hyperparameters for label-free training, following TTRL.
Hyperparameter Value
Train Batch Size 8
PPO Mini-Batch Size 1 (effective size of 32)
PPO Micro-Batch Size 2
Rollouts for Majority Vote 64
Rollouts Used for Training 32
Generation Temperature 1.0
Validation Temperature 0.6
Learning Rate 5e-7
Use KL Loss True
KL Loss Coefficient 0.001
Table 5: Key hyperparameters specific to the Evol-RL framework.
Hyperparameter Value
Asymmetric Clipping High ($\epsilon_{\text{high}}$) 0.28
Entropy Regularizer Coefficient ($\lambda_{\text{ent}}$) 0.003
Novelty Score Mixing Coefficient ($\alpha$) 0.5
A.6 Computational Resources
All experiments reported in this paper were conducted on a single server equipped with 8x NVIDIA H20 GPUs.
Appendix B Additional Experimental Results
Figure 5: Training dynamics of the majority-vote accuracy (maj@16) for Evol-RL and TTRL. Each panel plots the accuracy of the consensus answer derived from 16 rollouts over the course of training. The training datasets are: (Left) MATH-TRAIN, (Middle) MATH-500, and (Right) AIME24.
B.1 Analysis of the Majority Vote Signal
To further investigate the differences between Evol-RL and TTRL, we analyze the quality of the training signal itself by tracking the accuracy of the majority vote (maj@16) over the course of training, as shown in Figure 5. This analysis reveals how the self-generated pseudo-labels evolve under each method.
A highly consistent pattern emerges across all three training datasets. TTRL initially improves the maj@16 accuracy over the base model, but it quickly converges to a performance plateau. For the remainder of the training, its maj@16 accuracy remains largely unchanged, indicating that the consensus-only approach rapidly finds a local optimum for the consensus answer and becomes locked in, unable to discover better solutions.
In contrast, Evol-RL exhibits a markedly different dynamic. While its initial trajectory often mirrors that of TTRL, reflecting the early stabilizing influence of the consensus signal, a clear divergence occurs. Consistent with the inflection point observed in our main training dynamics analysis, Evol-RL’s maj@16 accuracy breaks away from the TTRL plateau and begins a second, sustained ascent. It reliably climbs to and stabilizes at a significantly higher level of accuracy. This demonstrates that Evol-RL’s exploration mechanisms not only improve the final policy but also progressively refine the quality of the pseudo-labels used for training, allowing the model to escape suboptimal consensus and continuously improve its understanding of the task.
B.2 Training Dynamics of 8B Models
Figure 6: Training dynamics for Evol-RL and TTRL on Qwen3-8B-Base model. Left: models trained on MATH-TRAIN. Middle: models trained on MATH-500. Right: models trained on AIME24. Each panel plots, over training steps, (i) Pass@1 on AIME25, (ii) average response length on the training set, and (iii) policy entropy on the training set.
The training dynamics of the 8B models, presented in Figure 6, largely mirror the patterns observed with the 4B models, confirming that the core mechanisms of EVOL-RL are robust to scale.
Across all three training datasets (MATH-TRAIN, MATH-500, and AIME24), we observe the same two-stage process. In Stage 1, both TTRL and EVOL-RL experience an initial drop in policy entropy and response length due to the strong initial pressure of the majority-vote signal. TTRL becomes permanently trapped in this low-entropy, low-complexity state.
In Stage 2, EVOL-RL consistently diverges at an "evolving point." Its policy entropy begins a sustained recovery, followed by a coordinated increase in average response length and out-of-domain accuracy on AIME25. This confirms that even at a larger scale, EVOL-RL successfully prevents entropy collapse and fosters a positive feedback loop where exploration, reasoning complexity, and performance reinforce one another, while the consensus-only TTRL approach stagnates.